# Sea ice floe segmentation in close-range optical imagery using active contour and foundation models

### v4

@@ Abstract
The size of sea ice floes in the marginal ice zone (MIZ) is a key factor influencing ice coverage, albedo, wave propagation through ice-covered waters, and ocean–atmosphere energy exchanges. Floe size can be observed by processing visual-range imagery from ships, aircraft, or satellites. However, autonomously capturing floe boundaries from imagery remains challenging, particularly due to the heterogeneity of sea ice, which impairs boundary definition and reduces image clarity. This study evaluates the accuracy of sea ice floe segmentation using the gradient vector flow (GVF) active contour method, the deep learning-based Segment Anything Model (SAM), and a hybrid approach combining GVF and SAM. These methods are evaluated on a representative subset of a large dataset of close-range, high-resolution sea ice imagery, collected from cameras aboard an icebreaker during an Antarctic winter expedition. Spanning a wide range of ice conditions and image clarity in the MIZ, the subset provides a rigorous test bed of segmentation approaches. Their performance is assessed in terms of floe detection accuracy, floe size distribution, and ice concentration, with results compared against a manually segmented benchmark. The findings indicate that SAM, when used in prompt-driven mode, offers the best balance between accuracy and computational efficiency. Its strong performance in estimating sea ice concentration and detecting floes, while maintaining close agreement with benchmark floe size distributions, makes it suitable for real-time applications and scalable analysis of large imagery datasets.
@@ 1 Introduction
Accurate observations of floe geometry, coverage, size, and dynamics are key to advancing and validating modern sea ice models (e.g., Bennetts et al., 2017; Roach et al., 2018; Wang et al., 2023; Buckley et al., 2024; Day et al., 2024), as these metrics strongly influence sea ice evolution and related climate feedbacks. Large scale observational datasets capturing spatial and temporal variability across the MIZ are necessary to constrain floe size distribution models, which are sensitive to assumptions about floe geometry and evolution (Bateson et al., 2022; Horvat, 2022). Floe characterization has long relied on high-resolution imagery combined with image processing techniques, and is now routinely extended using satellite, airborne, and shipborne sensors (e.g., Alberello et al., 2022; Alsharay et al., 2022; Herman et al., 2021; Lu et al., 2016; Paget et al., 2001; Parmiggiani et al., 2018; Sandru et al., 2020; Toyota et al., 2006; Zhang and Hughes, 2023; Zhang and Skjetne, 2015; Weissling et al., 2009; Hwang et al., 2017). Manual approaches, such as hand-drawn floe delineation, can be accurate, but they are painstaking and time-intensive, and do not scale to the number of images required for robust statistical analyses (Horvat, 2022). Classical automated methods, such as threshold segmentation, rely on user-defined values, which can introduce inconsistencies due to variations in sea ice conditions across the image (Heyn et al., 2017; Lu et al., 2008; Alberello et al., 2019). To address this limitation, more advanced techniques, including k-means clustering (Weissling et al., 2009; Lu et al., 2016), the gradient vector flow (GVF) snake algorithm (Zhang and Skjetne, 2015), and watershed segmentation (Dumas-Lefebvre and Dumont, 2023; Parmiggiani et al., 2018), have been applied to improve automation and robustness in sea ice analysis. Despite their advantages, these methods face challenges when applied to complex atmospheric and ice conditions that reduce the clarity of sea ice imagery. Although they perform well under optimal conditions, accuracy and consistency decline markedly in images with reduced clarity. Factors such as coarse resolution, cloud cover, fog, variable lighting, and low contrast between ice and water reduce image clarity. Additionally, snow cover and the presence of interstitial ice, thin ice forming in the spaces between floes, can obscure floe boundaries, creating the illusion of a more uniform surface or larger floes. These challenges complicate the identification of individual floes, particularly when they are densely packed, and make automated analysis of large image datasets problematic.
@@ 2 Image acquisition
Around 90,000 images were collected during the expedition, covering both day and night, with nighttime imaging supported by flashlights illuminating the sea ice (see sample images in Fig. 1c–e). Consequently, the database spans a wide range of ice and atmospheric conditions, with approximately 70% capturing complex sea ice scenarios. These situations are characterized by low light (Fig. 1e), densely packed and irregularly shaped floes, and diffuse boundaries, where floe edges are indistinct or visually merge due to snow cover, interstitial ice, or low contrast (Fig. 1d). This diversity offers a robust testbed for evaluating image processing methods
@@ 4.1 Active contours model (GVF)
The final result is obtained by integrating outputs from each phase of the analysis (see Fig. 4i). On average, the computational time is approximately 70 s per floe on a 2.3 GHz Intel Core i9 8 core processor, with no GPU acceleration. Computational time is reported per floe, rather than per image, to avoid bias from variability in the number of floes per image. A summary of computational times is reported in Tables 1 and 2.
@@ 4.2.1 Automatic mask generator (SAM-A)
Geometrical properties of floes and sea ice concentration are extracted from the final unified binary mask. The runtime for segmentation and floe analysis is approximately 0.8 s per floe.
@@ 4.2.2 Prompt-based (SAM-P)
The runtime is around 0.3 s per floe. Thus, SAM-P is more than two times faster than SAM-A, and it segments about 15% more floes. The prompt-based mode bypasses the model’s internal point selection process, speeding up segmentation. In contrast, the automatic mode uses a fixed number of points, which often misses some floes, leading to fewer segments.
@@ 4.3 A combination of SAM and GVF (SAM-GVF)
The final SAM-GVF output(Fig. 6i) is processed to extract floe properties and sea ice concentration. Segmentation and analysis had a runtime of approximately 4 s per floe.
@@ 5.1 Benchmark dataset
Reference data for the number and size of floes, as well as ice concentration, are subsequently derived. Concurrently, the same sample is segmented using the automated segmentation methods (GVF, SAM-A, SAM-P, and the combined SAM-GVF). Over 2,800 floes were identified across the 20 images, providing a solid foundation for benchmarking (Arbelàez et al., 2011; Martin et al., 2001).
5.2 Visual analysis of segmentation performance on sample images
@@
Two sample images representing contrasting ice morphology, one with distinct ice floes and well-defined edges, and another featuring diffuse boundaries due to closely packed or interconnected floes and interstitial ice, are used to assess the different methods against manual segmentation (Figures 7 and 8). This comparison illustrates how varying sea ice conditions influence segmentation outcomes and provides insight into model performances (within their respective constraints). These include: missing floes, where floes identified in the manual segmentation are not detected by the automated methods; misinterpreted floes, which are either partially segmented or merged with adjacent floes; and “hallucinations”, referring to false objects where patterns or textures are mistakenly identified as floes (Tables 1 and 2).
@@
Figure 7 and Table 1 present the results for the image with distinct floe edges from the benchmark dataset (original photo shown in Fig. 1c). The chosen image has sharp contrast, uniform lighting, and low noise level, reducing ambiguities. Figure 8 and Table 2 report results for the more challenging case (original photo in Fig. 1d), featuring reduced contrast, subtle lighting variations, and interstitial ice. These factors complicate the detection of floe boundaries, making the automated segmentation more prone to merging distinct floes, overlooking small or faint floes, and introducing false floes where weak edges or texture gradients are mistaken for true boundaries.
@@ 5.2.1 Distinct ice floes with defined edges
The manual segmentation identifies 196 floes with an average diameter of 3.9 m and an ice concentration of 57%. The GVF method misses 35 floes, misinterprets 26, and produces 15 hallucinations (Fig. 7c), resulting in approximately 40% of floes not being correctly identified. The missing floes have an average diameter of 2.4 m, representing about 4% of the total surface area. Most misinterpretations are due to partial segmentation of relatively large floes (average diameter of 3.4 m), with few instances of adjacent floes being merged. The hallucinations consist of smaller floes, with an average diameter of 1.6 m. Despite these errors, the GVF method appears to capture the general characteristics of the floe population. However, the relatively high number of missing floes and hallucinations introduces uncertainties in the ice concentration when compared to the manual benchmark.
Table 1: Segmentation results for the distinct edges image using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Distinct edges image
@@
Figure 7: Visual comparison between manual and automated segmentation methods for an image with distinct ice floes and defined edges from the benchmark dataset. The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
@@
SAM-GVF misses ten floes, misinterprets four, and produces seven hallucinations ($\approx$10% of floes). The missing floes have an average diameter of 2.6 m, covering around 1% of the surface area, similar to SAM-A and SAM-P. Misinterpreted floes have an average diameter of 3.4 m and include both partially segmented and merged floes. Hallucinations have an average diameter of 1.8 m and align with those seen in SAM-P.
@@ 5.2.2 Diffuse sea ice boundaries
As expected, segmentation performance declines under more challenging conditions. For the sample image with diffuse boundaries, the manual segmentation detects 250 floes with an average diameter of 3 m and an ice concentration of 55%. The performance of GVF significantly degrades—it misses 104 floes, nearly three times the number observed in the distinct edges image. Misinterpretations also increase, with 40 floes affected, while the 17 hallucinations is similar to the distinct edges case (Fig. 8g). In total, $\approx$70% of floes are not interpreted correctly, where the missed floes have an average diameter of 1.8 m and cover approximately 7% of the image area. The misinterpretations have an average diameter of 3 m and result mostly from partial segmentation of larger floes. The hallucinations have an average diameter of 1.1 m and do not substantially affect the overall sea ice characteristics. However, the increased number of missed and misinterpreted floes reduces the reliability of ice concentration estimates in the diffuse boundaries image.
Table 2: Segmentation results for the diffuse boundaries image using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Diffuse boundaries image
@@
SAM-GVF 58 1.24 m 2% 20 2.84 m 3% 12 1.46 m 1% 5 s
Figure 8: Visual comparison between manual and automated segmentation methods for an image with diffuse sea ice boundaries from the benchmark dataset. The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
@@
SAM-GVF shows moderate performance degradation, missing 58 floes, which is a substantial increase (almost sixfold) compared to the distinct edges image. Misinterpretations remain relatively low, with 20 floes affected, while hallucinations increase slightly to 12, which is fewer than SAM-P and GVF (Fig. 8g). Similar to SAM-P, $\approx$36% of the floes are not captured correctly. The missed floes have an average diameter of 1.2 m, covering about 2% of the image area. Misinterpretations, with an average floe diameter of 2.8 m, mainly result from partial segmentation of larger floes and those with mixed snow and bare ice surfaces, which slightly distorts the floe size distribution. Hallucinations remain small, with an average diameter of 1.5 m, and do not significantly impact the overall sea ice characteristics.

### v5

@@ Abstract
The size of sea ice floes in the marginal ice zone (MIZ) is a key factor influencing ice coverage, albedo, wave propagation, and ocean–atmosphere energy exchanges. Floe size can be observed by processing visual-range imagery from ships, aircraft, or satellites. However, autonomously capturing floe boundaries remains challenging, particularly due to sea ice heterogeneity, which impairs boundary definition and reduces image clarity. This study evaluates the accuracy of sea ice floe segmentation using the gradient vector flow (GVF) active contour method, the deep learning-based Segment Anything Model (SAM), and a hybrid approach combining GVF and SAM. Methods are evaluated on a representative subset of a large dataset of close-range, high-resolution imagery collected from cameras aboard an icebreaker during an Antarctic winter expedition. Spanning a wide range of ice conditions and image clarity in the MIZ, the subset provides a rigorous segmentation test bed. Performance is assessed in terms of floe detection accuracy, size distribution, and ice concentration, with results compared against a manually segmented benchmark. Results indicate SAM, in prompt-driven mode, offers the best balance between accuracy and computational efficiency. Its strong performance in estimating sea ice concentration and detecting floes, while maintaining close agreement with benchmark floe size distributions, makes it suitable for real-time applications and scalable analyses of large imagery datasets. Compared with SAM, the combined SAM-GVF method provides more accurate floe boundary delineation, although at much higher computational cost, and is therefore better suited for analyses requiring precise floe shapes.
@@ 1 Introduction
Accurate observations of floe geometry, coverage, size, and dynamics are key to advancing and validating modern sea ice models (e.g., Bennetts et al., 2017; Roach et al., 2018; Wang et al., 2023; Buckley et al., 2024; Day et al., 2024), as these metrics strongly influence sea ice evolution and related climate feedbacks. Large scale observational datasets capturing spatial and temporal variability across the MIZ are necessary to constrain floe size distribution models, which are sensitive to assumptions about floe geometry and evolution (Bateson et al., 2022; Horvat, 2022). Floe characterization has long relied on high-resolution imagery combined with image processing techniques, and is now routinely extended using satellite, airborne, and shipborne sensors (e.g., Alberello et al., 2022; Alsharay et al., 2022; Herman et al., 2021; Lu et al., 2016; Paget et al., 2001; Parmiggiani et al., 2018; Sandru et al., 2020; Toyota et al., 2006; Zhang and Hughes, 2023; Zhang and Skjetne, 2015; Weissling et al., 2009; Hwang et al., 2017). Manual approaches, such as hand-drawn floe delineation, can be accurate, but they are painstaking and time‑intensive, and do not scale to the number of images required for robust statistical analyses (Horvat, 2022). Classical automated methods, such as threshold segmentation, rely on user-defined values, which can introduce inconsistencies due to variations in sea ice conditions across the image (Heyn et al., 2017; Lu et al., 2008; Alberello et al., 2019). To address this limitation, more advanced techniques, including k-means clustering (Weissling et al., 2009; Lu et al., 2016), the gradient vector flow (GVF) snake algorithm (Zhang and Skjetne, 2015), and watershed segmentation (Dumas-Lefebvre and Dumont, 2023; Parmiggiani et al., 2018), have been applied to improve automation and robustness in sea ice analysis. Despite their advantages, these methods face challenges when applied to complex atmospheric and ice conditions that reduce the clarity of sea ice imagery. Although they perform well under optimal conditions, accuracy and consistency decline markedly in images with reduced clarity. Factors such as coarse resolution, cloud cover, fog, variable lighting, and low contrast between ice and water reduce image clarity. Additionally, snow cover and the presence of interstitial ice, thin ice forming in the spaces between floes, can obscure floe boundaries, creating the illusion of a more uniform surface or larger floes. These challenges complicate the identification of individual floes, particularly when they are densely packed, and make automated analysis of large image datasets problematic.
@@ 2 Image acquisition
Around 90,000 images were collected during the expedition, covering both day and night, with nighttime imaging supported by flashlights illuminating the sea ice (see sample images in Fig. 1c–e). Consequently, the database spans a wide range of ice and atmospheric conditions, with approximately 70% capturing complex sea ice scenarios. These situations are characterized by low light (Fig. 1e), densely packed and irregularly shaped floes, and diffuse boundaries, where floe edges are indistinct or visually merge due to snow cover, interstitial ice, or low contrast (Fig. 1d).
@@ 4.1 Active contours model (GVF)
The final result is obtained by integrating outputs from each phase of the analysis (see Fig. 4i). On average, the computational time ranges from approximately 35 s to 75 s per floe, depending on floe morphology (i.e., shape, edge clarity, and surface texture), on a 2.3 GHz Intel Core i9 8-core processor without GPU acceleration. Computational time is reported per floe, rather than per image, to avoid bias from variability in the number of floes per image. A summary of computational times is reported in Tables 2, 3, 4, and 5.
@@ 4.2.1 Automatic mask generator (SAM-A)
Geometrical properties of floes and sea ice concentration are extracted from the final unified binary mask. The runtime for segmentation and floe analysis is approximately 0.8 s per floe regardless of floe morphology.
@@ 4.2.2 Prompt-based (SAM-P)
The total runtime is around 0.3 s per floe, including the generation of seeds from the distance transform matrix, making SAM-P more than twice as fast as SAM-A. This is because the prompt-based mode bypasses SAM’s internal point selection process, speeding up segmentation.
@@ 4.3 A combination of SAM and GVF (SAM-GVF)
The final SAM-GVF output(Fig. 6i) is processed to extract floe properties and sea ice concentration. Segmentation and analysis had a runtime ranging from approximately 4 s to 6 s per floe.
@@ 5.1 Benchmark dataset
Reference data for the number and size of floes, as well as ice concentration, are subsequently derived. A total of 2,874 floes are identified across the twenty benchmark images, with an average diameter of 3.3 m and an average ice concentration of 45%, providing a solid foundation for benchmarking (Arbelàez et al., 2011; Martin et al., 2001). The same images are segmented using the automated methods (GVF, SAM-A, SAM-P, and the combined SAM-GVF). Segmentation performance of these methods is summarized in Table 1, categorized by three types of errors: missing floes, where floes identified in the manual segmentation are not detected by the automated methods; misinterpreted floes, which are either partially segmented or merged with adjacent floes; and “hallucinations”, referring to false objects where patterns or textures are mistakenly identified as floes.
5.2 Overall segmentation performance
Compared with manual segmentation of the 20-image dataset, the GVF method exhibits the highest number of missed floes (1,030), accounting for approximately 35% of the total. SAM-P, in contrast, misses the fewest floes ($\approx$20%), while SAM-A and SAM-GVF each miss slightly more than 25%. Across all methods, missed floes have an average diameter of around 2 m, covering between 2% and 5% of the benchmark image area. The exception is SAM-A, for which missed floes are slightly smaller (1.7 m), suggesting that SAM, when segmenting autonomously without seed guidance, preferentially detects larger floes while overlooking smaller ones.
Benchmark dataset
Method Missing Misinterpreted Hallucinations
GVF 1030 2.09 m 5% 423 3.54 m 7% 198 1.79 m 1%
SAM-A 767 1.66 m 2% 241 5.37 m 8% 97 3.40 m 1%
SAM-P 579 2.00 m 2% 229 4.45 m 5% 654 1.13 m 1%
SAM-GVF 707 1.97 m 3% 167 3.99 m 3% 384 1.36 m 1%
Misinterpretations range from 6% to 15% of the total floes, with SAM-GVF performing best (167 misinterpretations) and GVF performing worst (423 misinterpretations). These misinterpreted floes typically have average diameters ranging between 3.5 m and 5.4 m, larger than those of the missed floes, indicating that these errors mostly result from merging adjacent floes into larger objects.
Hallucinations vary significantly among methods but generally affect only around 1% of the total image area, which can be considered marginal in terms of their impact on overall segmentation accuracy. The average diameter of hallucinated floes is typically below 2 m, except for SAM-A, which produces hallucinations averaging 3.4 m. This further confirms that SAM-A, even when incorrectly segmenting, tends to detect relatively larger objects. The marked variability in hallucinations, particularly notable for SAM-P and SAM-GVF, primarily arises from image “22-07-24 02-52-21” (Fig. 13), which alone accounts for more than 35% of their total hallucinations. In this image, snowfall is misclassified as small floes, substantially increasing the hallucination count. For such cases, none of the methods may be optimal, and alternative, more stringent pre-processing strategies, such as cropping images to retain only areas of higher clarity, should be considered to effectively reduce noise and improve segmentation accuracy.
Overall, these results highlight notable variability in segmentation performance across different methods. A detailed analysis of specific cases is presented in the following section to further clarify how segmentation methods respond to the inherent heterogeneity in sea ice and image clarity.
5.3 Segmentation performance on sample images
@@
Four sample images representing contrasting ice conditions are used to evaluate the segmentation methods against manual segmentation, each highlighting specific challenges. Two cases primarily focus on floe boundaries: one features distinct floes with well-defined edges, while the other exhibits diffuse boundaries due to closely packed or interconnected floes and interstitial ice. The remaining two cases examine floe shape: one with uniformly round circular floes, and another characterized by angular square-shaped floes. This comparison illustrates how varying sea ice conditions influence segmentation outcomes and provides deeper insight into model performances (within their respective constraints).
@@
Figure 7 and Table 2 present results for the image with distinct floe edges (“22-07-23 12-06-06” in Fig. 13). This image has sharp contrast, uniform lighting, and low noise level, reducing ambiguities. Figure 8 and Table 3 report results for the more challenging diffuse boundaries case (“22-07-23 09-58-14” in Fig. 13), featuring reduced contrast, subtle lighting variations, and interstitial ice. These factors complicate the detection of floe boundaries, making the automated segmentation more prone to merging distinct floes, overlooking small or faint floes, and introducing false floes where weak edges or texture gradients are mistaken for true boundaries.
Figure 9 and Table 4 illustrate results for the round circular floes image (“22-07-24 11-17-13” in Fig. 13), characterized by medium-sized floes with uniform shapes and rounded corners. Figure 10 and Table 5 detail results for the angular square floes image (“22-07-21 12-48-26” in Fig. 12), which contains generally larger, irregularly shaped floes alongside smaller floes with complex morphology. These angular floes, typically formed by merging several smaller pancakes, have intricate internal textures that present additional segmentation challenges. In both cases, floes are well separated, isolating shape as the primary factor influencing segmentation performance.
@@ 5.3.1 Distinct ice floes with defined edges
The manual segmentation identifies 196 floes with an average diameter of 3.9 m and an ice concentration of 57%. The GVF method misses 35 floes, misinterprets 29, and produces 12 hallucinations (Fig. 7c), resulting in approximately 40% of floes not being correctly identified. The missing floes have an average diameter of 2.4 m, representing about 4% of the total surface area. Most misinterpretations are due to partial segmentation of relatively large floes (average diameter of 3.3 m), with few instances of adjacent floes being merged. The hallucinations consist of smaller floes, with an average diameter of 1.3 m. Despite these errors, the GVF method appears to capture the general characteristics of the floe population. However, the relatively high number of missing floes and hallucinations introduces uncertainties in the ice concentration when compared to the manual benchmark.
Table 2: Segmentation results for the distinct floe edges image (“22-07-23 12-06-06”) using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Distinct floe edges image
@@
Figure 7: Visual comparison between manual and automated segmentation methods for an image with distinct ice floes and defined edges from the benchmark dataset (“22-07-23 12-06-06” in Fig. 13). The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
@@
SAM-GVF misses ten floes, misinterprets four, and produces seven hallucinations ($\approx$11% of floes). The missing floes have an average diameter of 2.6 m, covering around 1% of the surface area, similar to SAM-A and SAM-P. Misinterpreted floes have an average diameter of 3.4 m and include both partially segmented and merged floes. Hallucinations have an average diameter of 1.8 m and align with those seen in SAM-P.
@@ 5.3.2 Diffuse sea ice boundaries
As expected, segmentation performance declines under more challenging conditions. For the sample image with diffuse boundaries, the manual segmentation detects 250 floes with an average diameter of 3 m and an ice concentration of 55%. The performance of GVF significantly degrades—it misses 104 floes, nearly three times the number observed in the distinct edges image. Misinterpretations also increase, with 40 floes affected, while the 17 hallucinations is similar to the distinct edges case (Fig. 8c). In total, $\approx$65% of floes are not interpreted correctly, where the missed floes have an average diameter of 1.8 m and cover approximately 7% of the image area. The misinterpretations have an average diameter of 3 m and result mostly from partial segmentation of larger floes. The hallucinations have an average diameter of 1.1 m and do not substantially affect the overall sea ice characteristics. However, the increased number of missed and misinterpreted floes reduces the reliability of ice concentration estimates in the diffuse boundaries image.
Table 3: Segmentation results for the diffuse floe boundaries image (“22-07-23 09-58-14”) using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Diffuse floe boundaries image
@@
SAM-GVF 58 1.24 m 2% 20 2.84 m 3% 11 1.24 m <1% 5 s
Figure 8: Visual comparison between manual and automated segmentation methods for an image with diffuse sea ice boundaries from the benchmark dataset (“22-07-23 09-58-14” in Fig. 13). The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
@@
SAM-GVF shows moderate performance degradation, missing 58 floes, which is a substantial increase (almost sixfold) compared to the distinct edges image. Misinterpretations remain relatively low, with 20 floes affected, while hallucinations increase slightly to 11, which is fewer than SAM-P and GVF (Fig. 8i). Similar to SAM-P, $\approx$36% of the floes are not captured correctly. The missed floes have an average diameter of 1.2 m, covering about 2% of the image area. Misinterpretations, with an average floe diameter of 2.8 m, mainly result from partial segmentation of larger floes and those with mixed snow and bare ice surfaces, which slightly distorts the floe size distribution. Hallucinations remain small, with an average diameter of 1.2 m, and do not significantly impact the overall sea ice characteristics.